• Sent from LBL to Cambridge on 5 October 2004.
  • Sent from Cambridge to CERN on 24 January 2005.
  • In use in SR1 Test Setup
  • PRM updated to v17f on 6 June 2007
  • Was in SR1 on 11/9/2012, but faulty - only STARTING LED comes on
  • PROM replaced, now OK (Nov 2012)
  • Installed in crate 0 slot 20 on 20/11/2012
  • Removed 13/11/2013, due to flashing ST2 and BOC present LED goes off at configure step
  • Shipped to LBL 26.11.2013
  • Returned 10.04.2014

-- JohnHill - 06 Jun 2007

Edit | Attach | Watch | Print version | History: r9 < r8 < r7 < r6 < r5 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r9 - 2014-04-10 - DaveRobinson
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback